F=2i-3j N and r=3i=2j m then torque

Grade:Upto college level

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Lalshekar

Torque= r x F

         = (3i + 2j) X (2i - 3j)

         = 6 i X i - 9 i x j + 4 j x i - 6 j x j

         = 0 - 9k + 4 (-k) - 0

         = -13k  Nm   Answer

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.




vikas askiitian expert
13 years ago

Torque=cross product of force and displacement vector

   T=r*F

     =(3i+2j)*(2i-3j)

     ={ 6i*i - 9i*j + 4j*i - 6j*j}              

     i*i=j*j=0     and i*j=k, j*i=-k

    Torque=-13k                 

   *represents cross product